# Towards Understanding Why FixMatch Generalizes Better Than Supervised Learning

### v1

@@ 4.2 Results on Feature Learning Process
$\Phi_{i,l}^{(t)}:=\sum\nolimits_{r\in[m]}[\langle w_{i,r}^{(t)},v_{i,l}\rangle]^{+},\qquad(i\in[k],l\in[2])$
as a feature learning indicator of feature $v_{i,l}$ in class $i$ ($i\in[k],l\in[2]$), since it denotes the total positive correlation score between the $l$-th feature $v_{i,l}$ of class $i$ and all the $m$ convolution kernels $w_{i,r}\ (r\in[m])$ at the $t$-th iteration. Larger $\Phi_{i,l}^{(t)}$ means the network can better capture the feature $v_{i,l}$, and thus can better use feature $v_{i,l}$ for classification. See more discussion in Appendix D.
Next, FixMatch uses a confidence threshold ($\tau$) to govern the usage of unsupervised loss as in Eq. (5), delineating its feature learning process into Phase I and II. For Phase I, the network relies primarily on supervised loss due to its inability to generate highly confident pseudo-labels. As training continues, the network learns partial feature and becomes better at predicting highly confident pseudo-labels for unlabeled data. This marks the transition to Phase II, where unsupervised loss begins to contribute, driven by consistency regularization between weakly and strongly augmented samples.
@@
Suppose Assumptions 2,3 holds. For sufficiently large $k$ and $m=\mathrm{polylog}(k)$, by setting $\eta\leq 1/\mathrm{poly}(k)$ and $\tau=1-\tilde{O}(1/s^{2})$, with probability at least $1-e^{-\Omega(\log^{2}k)}$:
(a) FixMatch. At the end of Phase I which continues for $T=\frac{\mathrm{poly}(k)}{\eta}$ iterations, we have
$\Phi_{i,l}^{(T)}\geq\Omega(\log k),\ \ \Phi_{i,3-l}^{(T)}\leq 1/\mathrm{polylog}(k),\ \ (\forall i\in[k],\exists l\in[2]).$ (8)
At the end of its Phase II which continues another $T^{\prime}=\frac{\mathrm{poly}(k)}{\eta}$ iterations, we have
$\Phi_{i,l}^{(T+T^{\prime})}\geq\Omega(\log k),\quad(\forall i\in[k],\forall l\in[2]).$ (9)
(b) Supervised Learning. After running $T\geq\frac{\mathrm{poly}(k)}{\eta}$ iterations, Eq. (8) always holds.
See its proof in Appendix F. Theorem 5(a) indicates that Phase I in FixMatch continues for $T={\mathrm{poly}(k)/\eta}$ iterations. During this phase, the network learns only one of the two semantic features per class. Specifically, in Eq. (8), for any class $i\in[k]$, there exists an index $l\in[2]$ so that the correlation score $\Phi_{i,l}^{(T)}$ exceeds $\Omega(\log k)$, showing feature $v_{i,l}$ is captured; and the score $\Phi_{i,3-l}^{(T)}$ remains low, indicating failure of learning $v_{i,3-l}$. Then we analyze classification performance when Eq. (8) holds.
@@
See its proof in Appendix D. According to Corollary 6, after Phase I, the network can correctly classify the multi-view data, since each contains two features and the network has learned one of them. However, for single-view samples that possess only one of the features, the network’s classification accuracy is only around 50% since it may not have learned the specific feature present in the sample. Then by running another $T^{\prime}$ iterations in Phase II, as shown in Theorem 5(a), FixMatch enables the network to capture both two features $v_{i,1}$ and $v_{i,2}$ in class $i\ (\forall i\in[k])$. As indicated in Eq. (9), all features have large correlation scores $\Phi_{i,l}^{(T+T^{\prime})}$ $(\forall i\in[k],\forall l\in[2])$. By Corollary 6, for all training and test data, the network trained by FixMatch can correctly classify them with high probability. This explains the good classification performance of FixMatch in Theorem 4.
@@
For Phase II of FixMatch, the reason for it to learn the semantics missed in Phase I is as follows. Having learned one feature per class in Phase I, the network is capable of generating highly confident pseudo-labels for weakly-augmented multi-view samples. As the confidence threshold $\tau\!=\!1-\tilde{O}({1}/{s^{2}})$ is close to 1 (e.g., $\tau\!=\!0.95$), it ensures the correctness of these pseudo-labels. Then, FixMatch uses these correct pseudo-labels to supervise the corresponding strongly-augmented samples via consistency regularization. As shown in Eq. (7), strong augmentation $\mathcal{A}(\cdot)$ randomly removes the learned features in unlabeled multi-view samples with probabilities $\pi_{1}\pi_{2}$ or $(1-\pi_{1})\pi_{2}$, effectively converting these samples into single-view data containing the unlearned feature. Given the large volume of unlabeled data as specified in Assumption 2, these transformed single-view samples are significant in their size. Accordingly, they dominate the unsupervised loss, since the rest samples containing the learned feature are already correctly classified by the network after Phase I and contribute minimally to the training loss. Consequently, the unsupervised loss enforces the network to learn the unlearned feature in Phase II.
@@
For supervised learning (SL), Theorem 5(b) shows that with high probability, SL can only learn one of the two features for each class. This result accords with Phase I in FixMatch. Then according to Corollary 6, SL can correctly classify multi-view data using the single learned feature, but achieves only about 50% test accuracy on single-view data due to the unlearned feature, aligning with Sec. 4.1. By comparison, FixMatch achieves nearly 100% test accuracy on multi-view and single-view data, as it learns both semantic features for each class.
@@
Comparison to Other SSL Analysis. This work differs from previous works from two key aspects. a) Our work provides the first analysis for FixMatch-like SSLs on CNNs. In contrast, many other works (He et al., 2022; Ţifrea et al., 2023) analyze over-simplified models, e.g., linear learning models, that differs substantially from the highly nonlinear and non-convex networks used in SSL. Some other works (Rigollet, 2007; Singh et al., 2008; Van Engelen & Hoos, 2020) view the model as a black-box function and do not reveal insights to model design. b) This work is also the first one to reveal the feature learning process of SSL, deepening the understanding to SSL and unveiling the intrinsic reasons of the superiority of SSL over its SL counterpart.
@@ 4.3 Semantic-Aware FixMatch
In Theorem 13 of Appendix A, we prove that SA-FixMatch enjoys the same good training and test accuracy in Theorem 4, but reduces the required number of unlabeled data samples $N_{c}$ in vanilla FixMatch to $N_{u}\!=\!\max\{\pi_{1}\pi_{2},(1-\pi_{1})\pi_{2}\}N_{c}$, where $N_{c}$ is given in Assumption 2. This data efficiency stems from SA-FixMatch’s use of SA-CutOut, which selectively removes the well-learned features, thereby compelling the network to focus on learning previously missed or unlearned features. Detailed theoretical discussions and proofs are presented in Appendix E, illustrating how SA-FixMatch outperforms vanilla FixMatch in terms of data efficiency and better test performance.
@@ 5 Experiments
To corroborate our theoretical results, we evaluate SL, FixMatch, and SA-FixMatch on several image classification benchmarks, including CIFAR-100 (Krizhevsky et al., 2009), STL-10 (Coates et al., 2011), Imagewoof (Howard & Gugger, 2020), and ImageNet (Deng et al., 2009). Following standard SSL evaluation protocols (Sohn et al., 2020; Zhang et al., 2021a; Wang et al., 2022a), we use WRN-28-8 (Zagoruyko & Komodakis, 2016) for CIFAR-100, WRN-37-2 (Zhou et al., 2020) for STL-10 and Imagewoof, and ResNet50 (He et al., 2016) for ImageNet. We also apply SA-CutOut to other FixMatch-like SSL methods and compare their performance with the original algorithms. All experiments are run three times, and we report the mean and standard deviation. Additional experimental details are provided in Appendix K.4 and K.5.
@@ 5.1 Classification Results
Here we evaluate SL, FixMatch, and SA-FixMatch by using different number of labeled data on CIFAR100, STL-10, Imagewoof, and ImageNet. Table 1 shows that on STL-10, FixMatch and SA-FixMatch improve SL by a significant 28%+ of test accuracy under all three settings. From Tables 1 and 2, one can also observe very similar big improvement on other datasets, e.g., 13%+ on CIFAR-100 and Imagewoof, and 6%+ on ImageNet. These results verify the superiority of SSL methods over conventional SL methods, and are consistent with our theoretical findings in Sec. 4.1.
Table 1: Comparison of Test accuracy (%) on CIFAR-100 and STL-10.
CIFAR-100 STL-10
Label Amount 400 2500 10000 40 250 1000
SL 11.45 $\pm$ 0.12 40.45 $\pm$ 0.50 63.77 $\pm$ 0.29 23.61 $\pm$ 1.62 38.83 $\pm$ 1.12 64.08 $\pm$ 0.47
FixMatch 55.16 $\pm$ 0.63 71.36 $\pm$ 0.44 77.25 $\pm$ 0.22 70.00 $\pm$ 4.02 88.73 $\pm$ 0.92 93.45 $\pm$ 0.19
SA-FixMatch 55.57 $\pm$ 0.43 72.12 $\pm$ 0.20 77.46 $\pm$ 0.16 71.81 $\pm$ 4.23 89.45 $\pm$ 1.19 94.04 $\pm$ 0.19
@@
Meanwhile, from Tables 1 and 2, we find that our proposed SA-FixMatch outperforms vanilla FixMatch. For example, on Imagewoof, SA-FixMatch has a $1.5\%+$ average test accuracy improvement on FixMatch; on ImageNet, SA-FixMatch has a $1.38\%$ higher average test accuracy than FixMatch; on CIFAR-100 and STL-10, SA-FixMatch also consistently outperforms FixMatch, albeit with a smaller performance gap. The reason behind this different improvement is that for samples in CIFAR-100 and STL-10, the semantic subject in the image occupies the majority of the image (see Appendix K.6). Accordingly, a random square mask in CutOut can well remove the semantic features with high probability, and thus has very similar masking effects as our SA-CutOut. Indeed, if we reduce the mask size in CutOut, random mask has less chance to well mask the semantic features and its performance degrades as observed in Table 4. For the Imagewoof and ImageNet datasets, most semantic subject only occupies less than a quarter of the image (see Appendix K.6). Therefore, a random square mask in CutOut only has small probability to remove the semantic features compared with SA-CutOut, and SA-FixMatch has much better test performance than FixMatch.
Table 2: Comparison of Test accuracy (%) on Imagewoof and ImageNet.
Imagewoof ImageNet
Label Amount 250 1000 2000 100K
SL 25.94 $\pm$ 1.54 42.04 $\pm$ 0.90 60.77 $\pm$ 1.04 44.62 $\pm$ 1.16
FixMatch 43.00 $\pm$ 1.46 64.91 $\pm$ 1.18 74.05 $\pm$ 0.15 50.80 $\pm$ 0.73
SA-FixMatch 46.73 $\pm$ 1.36 67.76 $\pm$ 1.29 75.62 $\pm$ 0.13 52.18 $\pm$ 0.32
@@
The superior test accuracy of SA-FixMatch compared to FixMatch is consistent with our theoretical analysis in Sec. 4.3. The reason behind this improvement is that to achieve good test performance in Theorem 4, the Phase II of SSL algorithms need to effectively remove well-learned features for enforcing the network to learn missed semantic features in Phase I. While FixMatch uses CutOut to randomly mask learned features in unlabeled data, SA-FixMatch always well masks the learned semantic feature because it adopts SA-CutOut as discussed in Sec. 4.3. Therefore, with a fixed number of images in unlabeled dataset, SA-FixMatch can more effectively use unlabeled data for feature learning, and thus achieve better test performance.
@@ 5.2 Semantic Feature Learning
Figure 2: Visualization of WRN-28-8 via Grad-CAM on CIFAR-100. For each group of three images, the left, middle and right one are the visualization of the models trained with SL, FixMatch, and SA-FixMatch, respectively.
@@ 5.3 SA-CutOut on FixMatch Variants
Table 4: Effect of (SA-)CutOut mask size to test accuracy (%) on CIFAR-100 with 400 labeled data.

### v2

@@ 4.2 Results on Feature Learning Process
$\Phi_{i,l}^{(t)}:=\sum\nolimits_{r\in[m]}[\langle w_{i,r}^{(t)},v_{i,l}\rangle]^{+},\quad i\in[k],\quad l\in[2]$
as an indicator of feature learning for $v_{i,l}$ in class $i$. It represents the total positive correlation between feature $v_{i,l}$ and all $m$ convolution kernels $w_{i,r}$ ($r\in[m]$) at iteration $t$. A larger $\Phi_{i,l}^{(t)}$ indicates better capture and utilization of $v_{i,l}$ for classification. See Appendix D for further discussion.
Next, FixMatch applies a confidence threshold ($\tau$) to regulate the unsupervised loss in Eq. (5), dividing its feature learning process into Phase I and Phase II. In Phase I, the network relies primarily on supervised loss, as it cannot yet generate confident pseudo-labels. As training progresses, the network learns partial features, improving its ability to predict confident pseudo-labels for unlabeled data. This transition marks the start of Phase II, where the unsupervised loss plays a larger role, driven by consistency regularization between weakly and strongly augmented samples.
@@
Suppose Assumptions 2, 3 hold. For sufficiently large $k$ and $m=\mathrm{polylog}(k)$, setting $\eta\leq 1/\mathrm{poly}(k)$ and $\tau=1-\tilde{O}(1/s^{2})$ ensures that, with probability at least $1-e^{-\Omega(\log^{2}k)}$:
(a) FixMatch. At the end of Phase I, which runs for $T_{1}=\mathrm{poly}(k)/\eta$ iterations,
$\Phi_{i,l}^{(T_{1})}\geq\Omega(\log k),\quad\Phi_{i,3-l}^{(T_{1})}\leq 1/\mathrm{polylog}(k),\quad\forall i\in[k],\exists l\in[2].$ (8)
After Phase II, which runs for another $T_{2}=\mathrm{poly}(k)/\eta$ iterations,
$\Phi_{i,l}^{(T_{1}+T_{2})}\geq\Omega(\log k),\quad\forall i\in[k],\forall l\in[2].$ (9)
(b) Supervised Learning. After $T\geq\mathrm{poly}(k)/\eta$ iterations, Eq. (8) always holds.
See its proof in Appendix F. Theorem 5(a) indicates that Phase I in FixMatch continues for $T_{1}={\mathrm{poly}(k)/\eta}$ iterations. During this phase, the network learns only one of the two semantic features per class. Specifically, in Eq. (8), for any class $i\in[k]$, there exists an index $l\in[2]$ so that the correlation score $\Phi_{i,l}^{(T_{1})}$ exceeds $\Omega(\log k)$, showing feature $v_{i,l}$ is captured; and the score $\Phi_{i,3-l}^{(T_{1})}$ remains low, indicating failure of learning $v_{i,3-l}$. Then we analyze classification performance when Eq. (8) holds.
@@
See its proof in Appendix D. According to Corollary 6, after Phase I, the network can correctly classify multi-view samples, as each contains two semantic features and the network learns at least one of them. However, for single-view samples, which contain only one semantic feature, the classification accuracy is around 50% since the network may not learn the specific feature present. Then by running another $T_{2}$ iterations in Phase II, Theorem 5(a) shows that FixMatch enables the network to capture both semantic features $v_{i,1}$ and $v_{i,2}$ for each class $i\in[k]$. As indicated by Eq. (9), all features achieve large correlation scores $\Phi_{i,l}^{(T_{1}+T_{2})}$ for all $i\in[k],l\in[2]$. Therefore, by Corollary 6, the network trained by FixMatch can correctly classify all training and test samples with high probability, explaining the strong generalization performance observed in Theorem 4.
@@
For Phase II of FixMatch, the reason for it to learn the semantic features missed in Phase I is as follows. Having learned one semantic feature per class in Phase I, the network is capable of generating highly confident pseudo-labels for weakly-augmented multi-view samples. As the confidence threshold $\tau\!=\!1-\tilde{O}({1}/{s^{2}})$ is close to 1 (e.g., $\tau\!=\!0.95$), it ensures the correctness of these pseudo-labels. Then, FixMatch uses these correct pseudo-labels to supervise the corresponding strongly-augmented samples via consistency regularization. As shown in Eq. (7), strong augmentation $\mathcal{A}(\cdot)$ randomly removes the learned features in unlabeled multi-view samples with probabilities $\pi_{1}\pi_{2}$ or $(1-\pi_{1})\pi_{2}$, effectively converting these samples into single-view data containing the unlearned feature. Given the large volume of unlabeled data as specified in Assumption 2, these transformed single-view samples are significant in their size. Accordingly, they dominate the unsupervised loss, since the rest samples containing the learned feature are already correctly classified by the network after Phase I and contribute minimally to the training loss. Consequently, the unsupervised loss enforces the network to learn the unlearned feature in Phase II.
@@
For SL, Theorem 5(b) shows that with high probability, SL learns only one of the two features for each class, consistent with Phase I in FixMatch. By Corollary 6, SL can correctly classify multi-view data using the single learned feature but achieves only about 50% test accuracy on single-view data due to the unlearned feature, aligning with Sec. 4.1. In contrast, FixMatch achieves nearly 100% test accuracy on both multi-view and single-view data, as it learns both semantic features for each class.
The key difference between FixMatch and SL lies in the additional unsupervised loss, which essentially serves as consistency regularization. Beyond SSL, consistency regularization also plays a crucial role in other learning paradigms. For example, in self-supervised learning, it promotes the acquisition of richer semantic features during pretraining. Our theoretical analysis offers valuable insights into these broader settings, and we leave the exploration as future work.
@@
Comparison to Other SSL Analysis. This work differs from previous works from two key aspects. (a) Our work provides the first analysis for FixMatch-like SSLs on CNNs. In contrast, many other works (He et al., 2022; Ţifrea et al., 2023) analyze over-simplified models, e.g., linear learning models, that differs substantially from the highly nonlinear and non-convex networks used in SSL. Some other works (Rigollet, 2007; Singh et al., 2008; Van Engelen & Hoos, 2020) view the model as a black-box function and do not reveal insights to model design. (b) This work is also the first one to reveal the feature learning process of SSL, deepening the understanding to SSL and unveiling the intrinsic reasons of the superiority of SSL over its SL counterpart.
@@ 4.3 Semantic-Aware FixMatch
In Theorem 13 of Appendix A, we prove that SA-FixMatch enjoys the same good training and test accuracy in Theorem 4, but reduces the required number of unlabeled data samples $N_{u}$ in vanilla FixMatch to $N_{c}\!=\!\max\{\pi_{1}\pi_{2},(1-\pi_{1})\pi_{2}\}N_{u}$, where $N_{u}$ is given in Assumption 2. This data efficiency stems from SA-FixMatch’s use of SA-CutOut, which selectively removes the well-learned features, thereby compelling the network to focus on learning previously missed or unlearned features. Detailed theoretical discussions and proofs are presented in Appendix E, illustrating how SA-FixMatch outperforms vanilla FixMatch in terms of data efficiency and better test performance.
Moreover, our analysis of SA-FixMatch remains valid even when the labeled and unlabeled datasets contain the same images, as SA-CutOut deterministically removes the learned features. This ensures the network continues to learn new semantic features in Phase II even with the same images as unlabeled data. Experiments in Sec. 5.4 further confirm that SA-FixMatch outperforms FixMatch and SL in this setting, validating the effectiveness of SA-CutOut in enhancing feature learning.
@@ 5 Experiments
To corroborate our theoretical results, we evaluate SL, FixMatch, and SA-FixMatch on CIFAR-100 (Krizhevsky et al., 2009), STL-10 (Coates et al., 2011), Imagewoof (Howard & Gugger, 2020), and ImageNet (Deng et al., 2009). Following standard SSL evaluation protocols (Sohn et al., 2020; Zhang et al., 2021a; Wang et al., 2022a), we use WRN-28-8 (Zagoruyko & Komodakis, 2016) for CIFAR-100, WRN-37-2 (Zhou et al., 2020) for STL-10 and Imagewoof, and ResNet-50 (He et al., 2016) for ImageNet. We also apply SA-CutOut to other FixMatch-like SSL methods and compare their performance against the originals. All experiments are repeated three times, reporting the mean and standard deviation. Further experimental details are provided in Appendix K.4 and K.5.
@@ 5.1 Classification Results
Here we evaluate the generalization performance of SL, FixMatch, and SA-FixMatch under varying amounts of labeled data. Following standard SSL benchmarks (Sohn et al., 2020; Zhang et al., 2021a; Chen et al., 2023), we use the entire training dataset as unlabeled dataset.
@@
Table 1 shows that on STL-10, FixMatch and SA-FixMatch outperform SL by over 28% in test accuracy across all settings. Similar substantial improvements are observed on other datasets, such as 13%+ on CIFAR-100 and Imagewoof, and 6%+ on ImageNet. These results highlight the superiority of SSL methods over conventional SL and align with our theoretical findings in Sec. 4.1.
@@
Meanwhile, Table 1 shows that SA-FixMatch outperforms vanilla FixMatch across all datasets. On Imagewoof, it improves average test accuracy by over 1.5%, while on ImageNet, it achieves a 1.38% gain. On CIFAR-100 and STL-10, SA-FixMatch also consistently surpasses FixMatch, though with a smaller margin. This variation arises because, in CIFAR-100 and STL-10, the semantic subject occupies most of the image (see Appendix K.6), allowing a random square mask in CutOut to effectively remove partial semantic features with high probability, producing similar masking effects as SA-CutOut. However, reducing the CutOut mask size lowers the likelihood of masking semantic features, leading to performance degradation (Table 4). In contrast, on Imagewoof and ImageNet, where the semantic subject occupies less than a quarter of the image (see Appendix K.6), a random square mask in CutOut is less likely to remove semantic features, making SA-CutOut significantly more effective and resulting in SA-FixMatch achieving much better test performance than FixMatch.
Table 1: Comparison of Test Accuracy (%) using the entire training dataset as unlabeled data.
Dataset CIFAR-100 STL-10 Imagewoof ImageNet
Label Amount 400 2500 10000 40 250 1000 250 1000 2000 100K
SL 11.45 $\pm$ 0.12 40.45 $\pm$ 0.50 63.77 $\pm$ 0.29 23.61 $\pm$ 1.62 38.83 $\pm$ 1.12 64.08 $\pm$ 0.47 25.94 $\pm$ 1.54 42.04 $\pm$ 0.90 60.77 $\pm$ 1.04 44.62 $\pm$ 1.16
FixMatch 55.16 $\pm$ 0.63 71.36 $\pm$ 0.44 77.25 $\pm$ 0.22 70.00 $\pm$ 4.02 88.73 $\pm$ 0.92 93.45 $\pm$ 0.19 43.00 $\pm$ 1.46 64.91 $\pm$ 1.18 74.05 $\pm$ 0.15 50.80 $\pm$ 0.73
SA-FixMatch 55.57 $\pm$ 0.43 72.12 $\pm$ 0.20 77.46 $\pm$ 0.16 71.81 $\pm$ 4.23 89.45 $\pm$ 1.19 94.04 $\pm$ 0.19 46.73 $\pm$ 1.36 67.76 $\pm$ 1.29 75.62 $\pm$ 0.13 52.18 $\pm$ 0.32
@@
The superior test accuracy of SA-FixMatch over FixMatch aligns with our theoretical analysis in Sec. 4.3. To achieve strong test performance in Theorem 4, Phase II of SSL must effectively remove well-learned features to enforce learning of missed semantic features from Phase I. While FixMatch relies on CutOut to randomly mask learned features, SA-FixMatch consistently masks them using SA-CutOut (Sec. 4.3). As a result, with a fixed unlabeled dataset size, SA-FixMatch utilizes unlabeled data more effectively for feature learning, leading to better test performance.
@@ 5.2 Semantic Feature Learning
Figure 2: Visualization of WRN-28-8 via Grad-CAM on CIFAR-100. Each group of three images corresponds to models trained with SL (left), FixMatch (middle), and SA-FixMatch (right).
@@ 5.3 SA-CutOut on FixMatch Variants
Table 3: Comparison of Test Accuracy (%) using the same training dataset for (SA-)FixMatch and SL.
Dataset STL-10 CIFAR-100 ImageNet
Data Amount 40 250 1000 400 2500 10000 100K
SL 19.93 44.06 67.29 9.87 40.98 63.48 41.82
FixMatch 38.88 64.70 79.15 18.58 47.20 67.94 43.34
SA-FixMatch 40.25 65.85 79.74 19.72 47.71 68.30 44.88
Table 4: Effect of (SA-)CutOut mask size on test accuracy (%) on CIFAR-100 with 400 labeled data.
@@ 5.4 Ablation Study
Same Training Dataset
We evaluate (SA-)FixMatch and SL under the same training dataset setting, as described in Sec. 4.3. As shown in Table 4, SA-FixMatch significantly outperforms SL, reaffirming the superiority of SSL over SL and further validating our theoretical insights in Sec. 4.3. Moreover, SA-FixMatch surpasses FixMatch, demonstrating the effectiveness of our proposed method.
Maske Size